Measurements were taken for the workstation that seals ball bearings; the data is:

Mean 118.45
Standard Dev 1.25
Lower Spec 117.96
Upper Spec 119.15

Determine the capability index. (Keep four decimal places in your answer)
Answer #1

Answer: 0.13

Cpk (Capability Index) = min((Upper Specification - Mean)/3σ , (Mean-Lower Specification)/3σ))

=min((119.15-118.45/3*1.25) , (118.45-117.96/3*1.25)

= min(0.187, 0.13066)

By considering the minimum, capability index is 0.1307